# On the Gap Structure of Generalized Stirling Numbers

### v1

@@ Abstract
Katugampola’s 2015 study of generalized fractional differential operators produced triangular arrays of integer coefficients $c_{n,k}^{(r)}$ indexed by fractional order $r$ and dimensions $n,k$, yet no combinatorial interpretation has been established for any fractional order. We provide the first such interpretation, establishing a dual contribution: a complete combinatorial interpretation for $r=1/2$ and $n\in\{1,2,3\}$, together with a rigorous proof that this interpretation cannot extend to $n\geq 4$ within the same framework. For $n\in\{1,2,3\}$, we show that $c_{n,k}^{(1/2)}$ counts binary sequences satisfying two conditions: containing at least one $B$ symbol and having gap $\leq 1$, where the gap measures the distance between the first and last occurrence of $B$. Each sequence is assigned a type $k$ via a parity-dependent formula involving the gap value, and complete enumeration verifies exact agreement with Katugampola’s coefficients. However, we prove via an obstruction theorem that the gap $\leq 1$ constraint inherently limits any such model to producing at most two distinct types per row, while Katugampola’s sequence requires at least three types for all $n\geq 4$. Our characterization is complete: the gap $\leq 1$ binary sequence interpretation correctly enumerates $c_{n,k}^{(1/2)}$ if and only if $n\in\{1,2,3\}$. This work transforms a computational observation into a rigorous impossibility theorem, providing essential guidance for future attempts at complete combinatorial interpretations of fractional calculus coefficients.
Keywords: generalized Stirling numbers, fractional calculus, binary sequences, combinatorial interpretation, obstruction theorem.
@@ 1 Introduction
Fractional calculus, the study of derivatives and integrals of arbitrary real or complex order, extends the classical theory of integer-order differentiation and integration. The subject traces its origins to a 1695 correspondence between Leibniz and L’Hôpital regarding the meaning of a derivative of order $\alpha=1/2$, and has since found applications across diverse fields including viscoelasticity, anomalous diffusion, control theory, and mathematical finance. A central object of study in fractional calculus is the expansion of generalized fractional differential operators, which naturally give rise to triangular arrays of integer coefficients analogous to classical combinatorial structures such as Stirling numbers and binomial coefficients. Understanding the combinatorial interpretation of these coefficient arrays—that is, identifying discrete structures whose enumeration produces the coefficients—remains an important problem connecting analysis, operator theory, and combinatorics.
In a 2015 study of Mellin transforms, Katugampola [1] introduced a family of generalized fractional integral and differential operators that unify the classical Riemann–Liouville and Hadamard fractional operators into a single framework parameterized by a real number $\rho>0$. Through systematic analysis of the Mellin transform properties of these operators, Katugampola derived triangular arrays of integer coefficients $c_{n,k}^{(r)}$ indexed by fractional order $r\in\mathbb{Q}^{+}$ and dimensions $n,k\in\mathbb{N}$. For the case $r=1/2$, these coefficients form the triangular array shown in Table 1, which corresponds to OEIS sequence A223168 [2]. The coefficients exhibit intriguing arithmetic patterns: the odd rows ($n=1,3,5,\ldots$) relate to generalized Laguerre polynomials, while the even rows display symmetric structure reminiscent of binomial coefficients. Despite this regularity, no combinatorial interpretation—that is, no description in terms of counting discrete structures—has been previously established for any fractional order $r$ in Katugampola’s framework.
@@
Our obstruction result provides essential guidance for future research. It eliminates an entire class of potential models from consideration: any successful combinatorial interpretation for $n\geq 4$ must either abandon the gap $\leq 1$ restriction (perhaps using gap $\leq k-1$ or gap $\leq\lfloor n/2\rfloor$), modify the type assignment mechanism, or introduce fundamentally different combinatorial structures. Beyond the specific case $r=1/2$, our methodology demonstrates a general approach combining positive constructions, computational boundary identification, and rigorous obstruction proofs—a template applicable to investigating other fractional orders $r\in\{1/3,1/4,1/5,\ldots\}$ in Katugampola’s framework (OEIS sequences A223169–A223172). The appearance of binary sequences with gap constraints in the combinatorial interpretation of fractional operator coefficients is intriguing and suggests deeper connections between discrete structures and fractional calculus. Understanding why the gap $\leq 1$ structure suffices for $n\leq 3$ but fails for $n\geq 4$ may illuminate fundamental relationships between combinatorial enumeration and the analytic structure of generalized fractional differential operators.
@@ 2.3 Valid Sequences
The “has BB” condition ensures that the gap function is well-defined. The gap $\leq 1$ restriction concentrates the $B$ symbols—either at a single position or two consecutive positions. Sequences with gap $\geq 2$ are “too spread out” and are excluded.
@@ 5 The Success of the Gap $\leq 1$ Structure
Proposition 5.1.
For all $n\geq 1$, $|V_{n}|=2n-1$.
Proof.
There are $n$ sequences with $\mathrm{gap}=0$ (exactly one $B$ at position $i$, for $i=1,\ldots,n$) and n1n-1 sequences with $\mathrm{gap}=1$ (two consecutive $B$’s starting at position $i$, for $i=1,\ldots,n-1$). Thus $|V_{n}|=n+(n-1)=2n-1$. ∎
@@ 7.1 Future Directions
Other fractional orders. Katugampola’s framework generates sequences for all rational orders $r\in\{1/2,1/3,1/4,1/5,\ldots\}$ (OEIS A223168–A223172 [2]). For $r=1/3$, does the natural alphabet expand to ternary sequences $\{R,B_{1},B_{2}\}$?
Connection to Laguerre polynomials. Katugampola notes that the odd rows of the $r=1/2$ triangle relate to generalized Laguerre polynomials $L_{n}^{(1/2)}(x)$. Tools from orthogonal polynomial theory, including continued fractions and three-term recurrences, may be relevant.

### v2

@@ Abstract
Katugampola’s study of generalized fractional differential operators produced triangular arrays of integer coefficients $c_{n,k}^{(r)}$, yet no combinatorial interpretation has been established for any fractional order. We provide the first such interpretation, establishing a dual contribution: a complete combinatorial interpretation for $r=1/2$ and $n\in\{1,2,3\}$, together with a rigorous proof that this interpretation cannot extend to $n\geq 4$. For $n\in\{1,2,3\}$, we show that $c_{n,k}^{(1/2)}$ counts binary sequences with at least one $B$ symbol and gap $\leq 1$ (distance between first and last $B$). Each sequence is assigned a type $k$ via a parity-dependent gap formula. However, we prove via an obstruction theorem that the gap $\leq 1$ constraint limits any such model to two distinct types per row, while Katugampola’s sequence requires at least three for all $n\geq 4$. This characterization is complete: the gap $\leq 1$ interpretation works if and only if $n\in\{1,2,3\}$. This transforms a computational observation into a rigorous obstruction theorem, guiding future combinatorial interpretations of fractional coefficients.
Keywords: generalized Stirling numbers, fractional calculus, binary sequences, combinatorial interpretation, fractional differential operators
@@ 1 Introduction
Fractional calculus, the study of derivatives and integrals of arbitrary real or complex order, extends the classical theory of integer-order differentiation and integration. The subject traces its origins to a 1695 correspondence between Leibniz and L’Hôpital regarding the meaning of a derivative of order $\alpha=1/2$, and has since found applications across diverse fields including viscoelasticity, anomalous diffusion, control theory, and mathematical finance. A central object of study in fractional calculus is the expansion of generalized fractional differential operators, which naturally give rise to triangular arrays of integer coefficients analogous to classical combinatorial structures such as Stirling numbers and binomial coefficients [1]. Understanding the combinatorial interpretation of these coefficient arrays—that is, identifying discrete structures whose enumeration produces the coefficients—remains an important problem connecting analysis, operator theory, and combinatorics.
In a 2015 study of Mellin transforms, Katugampola [2] introduced a family of generalized fractional integral and differential operators that unify the classical Riemann–Liouville and Hadamard fractional operators [3] into a single framework parameterized by a real number $\rho>0$. Through systematic analysis of the Mellin transform properties of these operators, Katugampola derived triangular arrays of integer coefficients $c_{n,k}^{(r)}$ indexed by fractional order $r\in\mathbb{Q}^{+}$ and dimensions $n,k\in\mathbb{N}$. For the case $r=1/2$, these coefficients form the triangular array shown in Table 1, which corresponds to OEIS sequence A223168 [4]. The coefficients exhibit intriguing arithmetic patterns: the odd rows ($n=1,3,5,\ldots$) relate to generalized Laguerre polynomials [6], while the even rows display symmetric structure reminiscent of binomial coefficients. Despite this regularity, no combinatorial interpretation—that is, no description in terms of counting discrete structures—has been previously established for any fractional order $r$ in Katugampola’s framework.
@@
Our obstruction result provides essential guidance for future research. It eliminates an entire class of potential models from consideration: any successful combinatorial interpretation for $n\geq 4$ must either abandon the gap $\leq 1$ restriction (perhaps using gap $\leq k-1$ or gap $\leq\lfloor n/2\rfloor$), modify the type assignment mechanism, or introduce fundamentally different combinatorial structures. Beyond the specific case $r=1/2$, our methodology demonstrates a general approach combining positive constructions, computational boundary identification, and rigorous obstruction proofs—a template applicable to investigating other fractional orders $r\in\{1/3,1/4,1/5,\ldots\}$ in Katugampola’s framework (OEIS sequences A223169–A223172 [5]). The appearance of binary sequences with gap constraints in the combinatorial interpretation of fractional operator coefficients is intriguing and suggests deeper connections between discrete structures and fractional calculus. Understanding why the gap $\leq 1$ structure suffices for $n\leq 3$ but fails for $n\geq 4$ may illuminate fundamental relationships between combinatorial enumeration and the analytic structure of generalized fractional differential operators.
@@ 2.3 Valid Sequences
The “has BB” condition ensures that the gap function is well-defined. The gap $\leq 1$ restriction concentrates the $B$ symbols—either at a single position (gap =0=0) or spanning at most two consecutive positions (gap =1=1). Sequences with gap $\geq 2$ are “too spread out” and are excluded.
@@ 5 The Success of the Gap $\leq 1$ Structure
This sequence of consecutive odd numbers suggests an underlying combinatorial structure that warrants further investigation. Indeed, one can verify that $|V_{n}|=2n-1$ for all $n\geq 1$, since there are $n$ sequences with gap =0=0 (exactly one $B$ at some position $i$) and n1n-1 sequences with gap =1=1 (two consecutive $B$’s starting at some position $i\leq n-1$).
@@ 7.1 Future Directions
Other fractional orders. Katugampola’s framework generates sequences for all rational orders $r\in\{1/2,1/3,1/4,1/5,\ldots\}$ (OEIS A223168–A223172 [4, 5]). For $r=1/3$, does the natural alphabet expand to ternary sequences $\{R,B_{1},B_{2}\}$?
Connection to Laguerre polynomials. Katugampola notes that the odd rows of the $r=1/2$ triangle relate to generalized Laguerre polynomials $L_{n}^{(1/2)}(x)$ [6]. Tools from orthogonal polynomial theory, including continued fractions and three-term recurrences, may be relevant.